# Interpretable Clustering: A Survey

### v1

@@ 4 Interpretable pre-clustering methods
In the study of interpretable clustering models, while our goal is to achieve more transparent models, it is equally important to carefully consider the features used as model inputs to produce interpretable results. Specifically, existing interpretable pre-clustering methods, which focus on the research conducted prior to clustering, can be approached from two perspectives: (1) feature extraction and (2) feature selection. Although these two issues have been extensively studied in the field of machine learning, they are rarely connected to interpretability, especially in terms of how to mine features that are more easily understood by humans for subsequent clustering tasks. Therefore, we have compiled a list of papers identified through our exhaustive search related to interpretable feature extraction or selection before clustering, which we elaborate on in the following two subsections.
4.1 Feature extraction
Interpretable pre-clustering methods from the perspective of feature extraction typically focus on complex data types, such as multivariate time series (MTS). The extraction of meaningful and informative features can lead to the development of simpler models that better capture significant characteristics within complex data, thus enhancing interpretability and facilitating better understanding.
In the field of multivariate time series, the system presented in [16] automatically extracts features from the signals, encompassing both intra-signal features, which characterize each signal independently, and inter-signal features, which evaluate relationships between signals using interpretable metrics. To select the most important features, the authors propose two methods: an unsupervised mode employing Principal Feature Analysis (PFA) and a semi-supervised mode incorporating user annotations on small dataset samples, significantly reducing the number of features without compromising accuracy. Salles et al. [17] leverages adaptive gating in NNs to dynamically select the most relevant features for each instance. Using a Gumbel-SoftMax technique to handle discrete choices and annealed mean-squared error regularization to encourage sparsity, the model identifies features that contribute most to predicting performance. These selected features are then used for clustering, enhancing the relevance and interpretability of the clusters.
Drawing on Gestalt theory, an interpretable band selection algorithm [18] is proposed in which hyperspectral imagery is considered as continuously varying points based on proximity and continuity principles. The model, constructed using similarity and invariance principles, extracts three bands from the hyperspectral image sequence to form a pseudo-color image, enhancing consistency within categories and differences between categories. RGB colors are categorized into ten types, and the differences between the three channels and the standard colors are minimized using Euclidean distance, allowing pseudo-color mapping of different bands and intuitively displaying target differences within specific spectral bands, aligning with principles of visual perception.
4.2 Feature selection
Another category of interpretable pre-clustering methods focuses on accurately selecting features with strong discriminative power for different data structures from a set of redundant and complex features prior to clustering. These methods can significantly enhances the interpretability of clustering models while maintaining their accuracy.
Svirsky et al. [19] propose to train self-supervised local gates to learn a sample-specific sparse gate vector for each input. The learned vectors are then used for reconstruction via an autoencoder. This approach provides instance-level explanations by representing each sample through a selected feature set, allowing the model to utilize fewer features for each instance while maintaining interpretability.
To address the lack of interpretability in clustering of patient clinical event logs, Balabaeva et al. [20] propose a method that extends the feature set with binary features. Using Bayesian inference, they identify specific features associated with the clustering structure and compare these with the features used by experts when describing the clusters. This approach significantly enhances the interpretation of clinical pathways clustering.
Effenberger et al. [21] select a set of useful features using a greedy approach. This involves considering one feature at a time, starting with the highest-weighted feature, and selecting it unless it is extremely rare, used in nearly all solutions, or too similar to already selected features. Jaccard coefficient is utilized to measure the similarity between two features, which is the ratio of the intersection to the union of the sets of solutions containing these features.
@@ 5 Interpretable in-clustering methods
Interpretable in-modeling clustering methods serve as a direct source of interpretability within the broader category of interpretable clustering approaches, embedding interpretability within the algorithmic process of clustering itself. This form of interpretability is typically treated as an optimizable interpretability objective combined with conventional clustering criteria (e.g., SSE as used in $k$-means). Some methods approach the interpretability goal by incorporating it jointly with conventional clustering criteria as a multi-objective optimization problem [22], while most simply consider it as an additional term related to certain structural parameters [23].
There are two typical scenarios (S1 and S2) where interpretable in-clustering methods could easily be confused with their corresponding pre- or post-clustering methods, depending on the stage at which interpretability quality is considered:
S1: Is input from third-party algorithms required? The interpretable models used in these in-clustering methods can either directly induce a clustering result (e.g., using decision-tree models that derive clusters via tree growth) or collaborate with various algorithms’ costs through joint optimization of objective functions. These methods do not rely on or attach to reference clustering results from third-party algorithms. Even if some methods use initial clustering results as input, they remain agnostic to how the clustering cost is defined [24]. The boundary between these methods and post-clustering methods (which aim to explain existing clustering results) can sometimes be blurred. If the clustering is driven by explainability rather than by fitting a given third-party algorithm’s results with an approximation guarantee, then the method is more aligned with interpretable in-clustering approaches.
To more clearly illustrate the distinction between in-clustering and post-clustering methods, we can consider the following example:
Illustrative references to S1: Although both [25] and [23] optimize a specific explainability measure for the decision tree structure within their algorithms, the former represents a post-clustering method, while the latter is an in-clustering method. The method in [25] assumes a fixed reference clustering and fits a decision tree to that clustering, while reference [23], as stated in the paper, allows for variations in the reference clustering to potentially discover more explainable clusters. Therefore, they differ in terms of when interpretability is considered during the process, with the clustering tree model being utilized at different stages of clustering. The key emphasis of interpretable in-clustering methods is their exploratory nature during the clustering stage. This keeps the clustering results open to potential modification throughout the algorithmic process. When the clustering is derived from a black-box algorithm, any subsequent interpretation may be viewed as post-hoc rationalization, potentially making it less reliable. Ideally, trustworthy clustering results are produced directly by the interpretable model [14], reducing reliance on third-party clustering algorithms and enhancing transparency and controllability within the process.
S2: Are the features in the dataset inherently interpretable? Interpretable in-clustering methods handle various forms of data and adjust according to the characteristics of the dataset’s features. For typical vector data, the features are usually interpretable [26]: (1) for numerical features, cut values can be applied to split the feature vector by determining whether the feature values are greater or less than a threshold, which is a common approach in decision-tree-based clustering; (2) for categorical features, values can similarly be interpreted based on whether they include or exclude a specific category. However, for data such as social and biological networks, which lack explicit features [27], interpretable community detection methods aim to find concise descriptive features for nodes [28]. For images, whose features may lack inherent interpretability (e.g., pixel matrices without clear structural meaning along any given dimension), discovering structural or interpretable features becomes more challenging. In tasks that involve images with semantic content, such as in the field of descriptive clustering [29], the focus shifts to identifying interpretable tags. In sum, to handle those complex data with uninterpretable features, there is often a need to incorporate deep learning techniques [30, 31]. For categorical sequential datasets, where each sample is a discrete sequence of variable length, some conventional sequence clustering methods require transforming the sequences into feature vectors. However, this transformation often leads to a loss of interpretability from the original sequence space. Dong et al. [32] argue that Discriminative Sequential Pattern Mining is necessary before building interpretable clustering methods.
Certain methods closely integrate the search for interpretable features with the clustering process itself, which can blur the boundaries between in-clustering and pre-clustering methods. Those methods often emphasize interpretability at the cluster level, rather than at the object/instance level. Here are some examples of such methods that clearly illustrate how the process of extracting interpretable features is integrated into the in-clustering stage:
Illustrative references to S2: Kim et al. [33] propose a generative approach to identify distinguishing dimensions in high-dimensional binary data clustering, facilitating data exploration and hypothesis generation. Their system embeds interpretability criteria into the model, using logic-based feature extraction to group dimensions into interpretable sets that differentiate clusters. Huang et al. [34] develop a deep clustering algorithm for feature selection within clusters. Using $K$-parallel auto-reconstructive learning, based on graph Laplacian theory, their model learns distinct feature subsets by exploring unknown feature associations and performing automatic feature weighting to minimize cluster-specific loss, enhancing both clustering performance and interpretability.
After clarifying these two scenarios where in-clustering methods can be confused with pre- or post-clustering methods in certain contexts, the following subsections will further review and identify key aspects that define the research area of interpretable in-clustering. The discussion will focus on how interpretability objectives are integrated into the clustering algorithmic process, with particular attention given to typical types of interpretable models.
@@
TABLE I: Summary of various interpretable in-clustering methods, each listing the representative reference and corresponding criteria.
@@ 5.1 Decision tree-based methods
The decision tree model is widely recognized as an interpretable model in machine learning and is commonly used for classification and regression tasks. Its interpretability stems from the recursive, hierarchical splitting of data based on feature values to generate intermediate results, with the final output is traceable through the feature values used in the splits. Instances are distributed to different leaf nodes (clusters) determined by specific splitting points according to certain criteria, following a clear, transparent path from the root node (representing the whole dataset) down through the branch nodes, which is easily understood by end users.
Early attempts to apply decision trees to clustering can be found in [41], where uniformly distributed synthetic data were introduced as auxiliary data to build a standard (supervised) decision tree. This approach aimed to maximize the separation between the original data and the synthetic data by modifying the standard splitting criterion, such as information gain. Although this method used binary splits, which are relatively easy to understand, the reliance on data generation introduced additional assumptions, making it difficult to claim that the splits were truly interpretable. In contrast, [42] developed an unsupervised decision tree directly based on the original features. The authors proposed four different measures for selecting the most appropriate feature and two algorithms for splitting data at each branch node. However, to select a candidate splitting point for calculating these measures, preliminary steps were required to divide the numerical feature domain into intervals. A simpler splitting criterion and a more intuitive algorithmic framework is presented in [35] with the introduction of CUBT, which was further extended to categorical data in [43]. CUBT adopts a general approach similar to CART, involving three steps: maximal tree construction, followed by pruning and merging to simplify the tree structure. This unsupervised decision tree-based clustering model was also extended to the interpretable fuzzy clustering domain in [44], where fuzzy splitting at branch nodes was used to grow the initial tree, followed by merging similar clusters to create a more compact tree structure.
The aforementioned unsupervised decision tree-based models adopt a top-down approach, where all possible candidate splitting points are considered at the current branch node level, and criteria such as heterogeneity are calculated so that the tree grows greedily (greedy search) based on the optimal splits passed down from the parent node. However, this type of algorithm lacks global guidance, meaning that each split is optimized locally rather than achieving a globally optimized solution across the entire dataset.
@@ 5.2 Rule-based methods
The process of mining an optimal rule set to derive a specific cluster is often inspired by the field of pattern mining [50]. To ensure that different rule sets effectively correspond to their respective clusters, the rule set typically exhibits two key characteristics [51]: (1) frequency (meaningful), indicating that the rule set should cover as many samples within its corresponding cluster (true positives) as possible, and (2) discriminative power (unique), meaning that the rule set should minimize the number of samples mistakenly covered from other clusters (false positives).
To obtain a rule set for the purpose of interpretable clustering, a common approach is to start by quantifying interpretability based on how well a rule covers a specific cluster. For example, as demonstrated in [37], an interpretability score is defined to assess a feature value’s relevance to a cluster by considering the fraction of samples within the cluster that share that feature value. Given all candidate rules or rule sets (e.g., generated using frequent pattern mining), these methods aim to derive clusters that maximize the interpretability score while simultaneously optimizing cluster quality. Since interpretability objectives often conflict with cluster quality, existing methods typically incorporate the interpretability score as a user-specified bound to balance interpretability and cluster quality, alongside standard clustering objectives. The method in [22] introduces two explainability criteria for each rule set associated with a cluster: one similar to [37], and another that considers the distinctiveness of the rule set, meaning how few samples it covers outside of the associated cluster. Optimizing these two explainability objectives, together with cluster quality measures, is formulated into a multi-objective Mixed-Integer linear Optimization problem (multi-MIO). Furthermore, the method in [22] considers the maximum rule set length (lenRule), i.e., the number of feature values in the combination, as a constraint, ensuring that the created clusters are more interpretable by being represented through concise rules.
@@ 5.3 Other methods
In addition to the two widely used interpretable models mentioned above, other interpretable in-clustering methods create clusters or determine cluster membership based on representative elements, which can generally be categorized as boundary-based or centroid-like approaches. However, for these representative elements to be interpretable, certain properties need to be maintained. The following is a brief overview of these approaches.
Prototype (exemplar): In datasets where the original features are non-interpretable and difficult to understand, such as with images and text, especially when deep embeddings are used, recent work on interpretable in-clustering via exemplars has found that seeking high-level centroids can be useful for characterizing clusters and facilitating visualization. For example, [40] tackles the challenging problem of finding the minimum number of exemplars (nExemplar) without prior specification. Additionally, [31] proposes a new end-to-end framework designed to enhance scalability for larger datasets, making exemplar-based clustering more practical for real-world applications.
5.4 Summary
Various interpretable models, with others potentially existing and requiring further investigation, have been developed for in-clustering methods (summarized in Table I). These models consistently treat interpretability as a first-class objective, on par with clustering quality, incorporating it as an optimization target either directly or indirectly, depending on the model type. For instance, tree-based models often prioritize reducing the number of branch or leaf nodes, rule-based models focus on shorter rules, and geometric representation models, such as prototype-based models, aim to minimize the number of exemplars. More refined structural parameters as optimization targets require further research. For example, in literature [25], tree depth is considered an optimization target; however, this approach, designed to explain a given reference clustering result, belongs to post-clustering methods.
TABLE II: Summary of various interpretable post-clustering methods, each listing the representative reference and corresponding criteria.
@@ 5.4 Summary
6 Interpretable post-clustering methods
6.1 Decision tree-based methods
Research on the quality of decison tree constructed by interpretable post-clustering methods began with the work of Moshkovitz et al. [54]. They develop decision trees using a greedy approach that aims to minimize the number of errors at each split (i.e., the number of points separated from their corresponding reference cluster centers), stopping when the tree reaches $k$ leaf nodes. This method achieves an $O(k)$ approximation for the optimal $k$-medians and an $O(k^{2})$ approximation to the optimal $k$-means. Laber et al. [58] improve the approximation, achieving an $O(d\log k)$ approximation for optimal $k$-medians and an $O(kd\log k)$ approximation for the optimal $k$-means. They accomplish this by firstly construct $d$ decision trees, where $d$ is the number of dimensions in the data, then utilize these trees to build the final decision tree. The feature for splitting a node within the final decision tree is chosen based on the dimension with the maximum range among the centers contained in the current node. The specific feature value is associated with the node in the corresponding dimension’s decision tree, which is the least common ancestor (LCA) of the set of reference centers that reach the current node. Makarychev et al. [59] take a different approach by choosing splitting features and values that differentiate centers with greater distances within each node in a relatively random manner. This results in an $O(\log k\log\log k)$ approximation for the optimal $k$-medians and an $O(k\log k\log\log k)$ approximation for the optimal $k$-means. In the decision tree constructed in [60], the choice of cuts at each split node is entirely random, as long as it can separate different reference centers into different child nodes. It has been proven that this method can achieve an $O(\log^{2}k)$ approximation for the optimal $k$-medians and an $O(k\log^{2}k)$ approximation for the optimal $k$-means. Recentlty, Esfandiari et al. [61] focus on determining the maximum and minimum values of the reference centers along each dimension, sorting these values, and then sampling a split point that effectively separates the reference centers. Their method achieves an $O(\log k\log\log k)$ approximation for the optimal $k$-medians and an $O(k\log k)$ approximation for $k$-means. Several methods have been proposed to independently provide near-optimal algorithms for $k$-means or $k$-medians [62, 63, 64], which will not be elaborated upon here.
6.2 Rule-based methods
De Weerdt et al. [67] investigate the search for explanations for event logs by first generating feature sets from the data and then applying a best-first search procedure with pruning to construct the set of explanations. Through an iterative process, they continuously enhance the accuracy and conciseness of the explanations for the instances. Building on this work, Koninck et al. [68] mine concise rules for each individual instance from a black box support vector machine (SVM) model and discuss and evaluate different alternative feature sets that can be used as inputs for explanatory techniques.
6.3 Other methods

### v2

@@ 5 Interpretable pre-clustering methods
While most research on interpretable clustering focuses on generating transparent clustering outcomes, the interpretability of input features also plays a crucial role in ensuring understandable and trustworthy results. Existing interpretable pre-clustering methods aim to enhance feature interpretability at the data preparation stage and can be broadly categorized into two complementary perspectives: (1) interpretable feature extraction, which derives human-understandable representations from complex data, and (2) interpretable feature selection, which identifies compact and semantically meaningful subsets of features. Although both topics have been widely studied in the field of machine learning, they have rarely been systematically investigated in connection with subsequent clustering tasks.
From the perspective of feature extraction, studies have sought to obtain informative and human-understandable representations from complex data before clustering. Bonifati et al. [31] proposed Time2Feat to extract intra- and inter-signal features from multivariate time series using interpretable metrics and dimension reduction via Principal Feature Analysis or user annotation. Salles et al. [32] employed adaptive gating with Gumbel-SoftMax sampling to identify instance-relevant features that guide clustering, while Kang et al. [33] introduced an interpretable hyperspectral band-selection algorithm based on Gestalt principles, aligning extracted representations with human visual perception to enhance the transparency of clustering outcomes.
From the perspective of feature selection, methods focus on identifying compact and discriminative feature subsets that preserve clustering accuracy while improving interpretability. Svirsky et al. [34] trained self-supervised local gates to produce instance-specific sparse feature sets, revealing which attributes drive each cluster assignment. Effenberger et al. [35] applied a greedy feature selection strategy based on occurrence frequency and Jaccard similarity, generating concise and meaningful feature subsets that clarify the logic behind cluster formation.
@@ 6 Interpretable in-clustering methods
Interpretable in-clustering methods constitute the core of interpretable clustering research, as they embed interpretability directly within the clustering process rather than applying it before or after clustering. In such methods, interpretability is treated as an explicit optimization objective jointly considered with traditional clustering criteria such as the within-cluster sum of squared errors (SSE). Some approaches formulate it as a multi-objective optimization [36], while others impose it as a structural regularization term [37].
Clarifying methodological distinctions. In-clustering methods are sometimes confused with pre- or post-clustering approaches because of when interpretability is introduced. Two criteria help clarify their methodological scope:
  • Dependence on third-party algorithms. This distinguishes in-clustering methods from post-clustering ones. In-clustering integrates interpretability within the clustering objective, either by inducing clusters through interpretable models (e.g., tree-based clustering) or by jointly optimizing interpretability with standard costs, without relying on externally generated cluster labels [38]. In contrast, post-clustering methods explain reference partitions obtained from black-box algorithms. For example, both [39] and [37] optimize tree-based interpretability, yet the former fits a tree to fixed clustering results (post-clustering), whereas the latter jointly optimizes clustering and interpretability (in-clustering). Thus, in-clustering emphasizes exploratory clustering guided by interpretability, producing clusters directly through transparent models [21].
  • Interpretability during the clustering process. This criterion distinguishes in-clustering from pre-clustering methods. Pre-clustering enhances feature interpretability before clustering, whereas in-clustering jointly learns interpretable features and cluster structures. For tabular data [40], interpretability is straightforward through thresholds or categorical inclusion. For more complex data such as networks [41, 42], images, and sequences, features lack clear semantics; thus, in-clustering methods aim to discover clusters and meaningful representations simultaneously. In networks, concise node descriptors are extracted for interpretable community detection [43]; in images, semantic tags are identified via descriptive clustering [44]; and deep models improve interpretability through latent representation learning [45, 46]. For sequential data, vectorization often obscures meaning, motivating discriminative sequential pattern mining [47]. Some models integrate feature interpretability directly into clustering objectives: Kim et al. [48] group binary dimensions into logic-based interpretable sets, while Huang et al. [49] jointly optimize feature selection and clustering via a deep $K$-parallel auto-reconstructive framework with graph Laplacian regularization.
After clarifying these distinctions, the following subsections review representative in-clustering approaches, focusing on how interpretability objectives are embedded within clustering algorithms and realized through various interpretable models.
6.1 Decision tree-based methods
The decision tree model is widely recognized as an interpretable model in machine learning [50] and is commonly used for classification and regression tasks. Its interpretability stems from the recursive, hierarchical splitting of data based on feature values to generate intermediate results, and the final output is traceable through the feature values used in the splits. Instances are distributed to different leaf nodes (clusters) determined by specific splitting points according to certain criteria, following a clear, transparent path from the root node (representing the whole dataset) down through the branch nodes, which is easily understood by end users.
Early attempts to apply decision trees to clustering can be found in [51], where uniformly distributed synthetic data were introduced as auxiliary data to build a standard (supervised) decision tree. This approach aimed to maximize the separation between the original data and the synthetic data by modifying the standard splitting criterion, such as information gain. Although this method used binary splits, which are relatively easy to understand, the reliance on data generation introduced additional assumptions, making it difficult to claim that the splits were truly interpretable. In contrast, [52] developed an unsupervised decision tree directly based on the original features. The authors proposed four different measures for selecting the most appropriate feature and two algorithms for splitting data at each branch node. However, to select a candidate splitting point for calculating these measures, preliminary steps were required to divide the numerical feature domain into intervals [53, 54]. A simpler splitting criterion and a more intuitive algorithmic framework is presented in [55] with the introduction of CUBT, which was further extended to categorical data in [56]. CUBT adopts a general approach similar to CART [57], involving three steps: maximal tree construction, followed by pruning and merging to simplify the tree structure. This unsupervised decision tree-based clustering model was also extended to the interpretable fuzzy clustering domain in [58], where fuzzy splitting at branch nodes was used to grow the initial tree, followed by merging similar clusters to create a more compact tree structure.
The aforementioned unsupervised decision tree-based models adopt a top-down approach [59, 60], where all possible candidate splitting points are considered at the current branch node level, and criteria such as heterogeneity are calculated so that the tree grows greedily (greedy search) based on the optimal splits passed down from the parent node. However, this type of algorithm lacks global guidance, meaning that each split is optimized locally rather than achieving a globally optimized solution across the entire dataset.
6.2 Rule-based methods
The process of mining an optimal rule set to derive a specific cluster is often inspired by the field of pattern mining [67]. To ensure that different rule sets effectively correspond to their respective clusters, the rule set typically exhibits two key characteristics [68]: (1) frequency (meaningful), indicating that the rule set should cover as many samples within its corresponding cluster (true positives) as possible, and (2) discriminative power (unique), meaning that the rule set should minimize the number of samples mistakenly covered by other clusters (false positives).
To obtain a rule set for the purpose of interpretable clustering, a common approach is to start by quantifying interpretability based on how well a rule covers a specific cluster. For example, as demonstrated in [69], an interpretability score is defined to assess a feature value’s relevance to a cluster by considering the fraction of samples within the cluster that share that feature value. Given all candidate rules or rule sets (e.g., generated using frequent pattern mining), these methods aim to derive clusters that maximize the interpretability score while simultaneously optimizing cluster quality. Since interpretability objectives often conflict with cluster quality, existing methods typically incorporate the interpretability score as a user-specified bound to balance interpretability and cluster quality, alongside standard clustering objectives. The method in [36] introduces two explainability criteria for each rule set associated with a cluster: one similar to [69], and another that considers the distinctiveness of the rule set, meaning how few samples it covers outside the associated cluster. Optimizing these two explainability objectives, together with cluster quality measures, is formulated into a multi-objective Mixed-Integer linear Optimization problem (multi-MIO). Furthermore, the method in [36] considers the maximum rule set length (lenRule), i.e., the number of feature values in the combination, as a constraint, ensuring that the created clusters are more interpretable by being represented through concise rules.
6.3 Other methods
In addition to the two widely used interpretable models mentioned above, other interpretable in-clustering methods create clusters or determine cluster membership based on representative elements [72, 73], which can generally be categorized as boundary-based or centroid-like approaches. However, for these representative elements to be interpretable, certain properties need to be maintained. The following is a brief overview of these approaches.
Prototype (exemplar): In datasets where the original features are non-interpretable and difficult to understand, such as with images [76] and text, especially when deep embeddings are used, recent work on interpretable in-clustering via exemplars has found that seeking high-level centroids can be useful for characterizing clusters and facilitating visualization. For example, [77] tackles the challenging problem of finding the minimum number of exemplars (nExemplar) without prior specification. Additionally, [46] proposes a new end-to-end framework designed to enhance scalability for larger datasets, making exemplar-based clustering more practical for real-world applications.
@@ 6.3 Other methods
TABLE II: Summary of various interpretable in-clustering methods, each listing the representative reference and corresponding criteria.
@@
6.4 Summary
Various interpretable models have been developed for in-clustering methods, as summarized in Table II. Currently, most existing approaches in the literature either build upon these representative methods or follow similar principles that can be subsumed under them. These models consistently treat interpretability as a first-class objective, on par with clustering quality, incorporating it as an optimization target either directly or indirectly, depending on the model type. For instance, tree-based models often prioritize reducing the number of branch or leaf nodes, rule-based models focus on shorter rules, and geometric representation models, such as prototype-based models, aim to minimize the number of exemplars. More refined structural parameters as optimization targets require further research. For example, in literature [39], tree depth is considered an optimization target; however, this approach, designed to explain a given reference clustering result, belongs to post-clustering methods.
7 Interpretable post-clustering methods
@@ 7.1 Decision tree-based methods
Research on the quality of decison tree constructed by interpretable post-clustering methods began with the work of Moshkovitz et al. [78]. They develop decision trees using a greedy approach that aims to minimize the number of errors at each split (i.e., the number of points separated from their corresponding reference cluster centers), stopping when the tree reaches $k$ leaf nodes. This method achieves an $O(k)$ approximation for the optimal $k$-medians and an $O(k^{2})$ approximation to the optimal $k$-means. Laber et al. [79] improve the approximation, achieving an $O(d\log k)$ approximation for optimal $k$-medians and an $O(kd\log k)$ approximation for the optimal $k$-means. They accomplish this by firstly constructing $d$ decision trees, where $d$ is the number of dimensions in the data, then utilize these trees to build the final decision tree. The feature for splitting a node within the final decision tree is chosen based on the dimension with the maximum range among the centers contained in the current node. The specific feature value is associated with the node in the corresponding dimension’s decision tree, which is the least common ancestor (LCA) of the set of reference centers that reach the current node. Makarychev et al. [80] take a different approach by choosing splitting features and values that differentiate centers with greater distances within each node in a relatively random manner. This results in an $O(\log k\log\log k)$ approximation for the optimal $k$-medians and an $O(k\log k\log\log k)$ approximation for the optimal $k$-means, with $\log\log k$ denoting the iterated logarithm $\log(\log k)$. In the decision tree constructed in [81], the choice of cuts at each split node is entirely random, as long as it can separate different reference centers into different child nodes. It has been proven that this method can achieve an $O(\log^{2}k)$ approximation for the optimal $k$-medians and an $O(k\log^{2}k)$ approximation for the optimal $k$-means. Recentlty, Esfandiari et al. [82] focus on determining the maximum and minimum values of the reference centers along each dimension, sorting these values, and then sampling a split point that effectively separates the reference centers. Their method achieves an $O(\log k\log\log k)$ approximation for the optimal $k$-medians and an $O(k\log k)$ approximation for $k$-means. Several methods have been proposed to independently provide near-optimal algorithms for $k$-means or $k$-medians [83, 84, 85], which will not be elaborated upon here.
@@ 7.2 Rule-based methods
De Weerdt et al. [88] investigate the search for explanations for event logs by first generating feature sets from the data and then applying a best-first search procedure with pruning to construct the set of explanations. Through an iterative process, they continuously enhance the accuracy and conciseness of the explanations for the instances. Building on this work, Koninck et al. [89] mine concise rules for each individual instance from a black-box support vector machine (SVM) model and discuss and evaluate different alternative feature sets that can be used as inputs for explanatory techniques.
Ofek et al. [90] introduced Cluster-Explorer, a post-hoc framework that explains black-box clustering pipelines through compact rule sets. The method formulates explanation generation as a multi-objective optimization over coverage, separation error, and conciseness, which is efficiently solved using a generalized frequent itemset mining (gFIM)-based search. By representing predicates as items and applying Pareto pruning, the algorithm derives minimal-length rule sets that maximize cluster coverage while minimizing overlap across clusters.
@@ 7.3 Other methods
TABLE III: Summary of various interpretable post-clustering methods, each listing the representative reference and corresponding criteria.